a school has 1200 seniors. 60% are girls the rest are boys. 70% of the girls in class vote for a prom and 40% of boys vote for a prom. what percent of students did not want the prom.
Answer by randyjoseph7(13)

1200 seniors
x.60 are girls=
720 girls which leaves 480 boys
70& of the girls would be 504 girls who voted yes
40% of boys would be 192 boys who voted yes
since 70% of the girls did that leaves 30% of girls vote no, 216 girls
and since 40% of boys did, that leaves 60% of boys who vote no,288 boys
216+288=504 students combined who dont want prom
so out of 1200 students, 504 vote no and 696 vote yes
cross multiply
1200/504 = 100/x
1200 nstudents 504 who domt want prom and 100% of school would be to 1200 students so you need to find out the percent who dont
turns out 504x100=50400 divided by 1200= 42%